OpenStudy (anonymous):

The point (-1,2) lies on a circle whose center is at (-4,1) . Find an equation of this circle.

OpenStudy (anonymous):

well you the know the equation of a circle is x^2 + y^2 = r^2 where x and y can translated on a graph and r is the radius. you know the radius is going to be 3 because of the distance between the x coordinates. so you have r. all that's left is to translate the x and y to be where the center is at. working from the origin (0,0) the circle is -4 in the x direction and +1 in the y direction. so correct me if i'm wrong somebody but the equation is (x+4)^2 + (y+1)^2 = 9 because i think the x is supposed to be opposite of how it's translated? or is it the y? but the equation should be very similar to that

OpenStudy (anonymous):

i could have sworn i just answered this

OpenStudy (anonymous):

\[(x + 4)^{2} + (y - 1)^{2} = 10\]

OpenStudy (anonymous):

and that is the answer i got.

OpenStudy (anonymous):

5x+4y=-7 y=2x+8 Use substitution to solve the system. I got x=-3 y=2